A Merchant mixes tea that sells for $7.20/kg and tea that sells for $9.20/kg to make 100 kg of tea that he sells for $8.80/kg. How many kilograms of each type of tea did he mix?
Also to solve it, it has to be either the substitution or elimination method.

Answer by checkley77(7072)

X+Y=100 OR X=100-Y
7.20X+9.20Y=8.80*100 NOW REPLACE X IN THISD EQUATION BY (100-Y) & SOLVE FOR X.
7.20(100-Y)+9.20Y=880
720-7.20Y+9.20Y=880
2Y=880-720
2Y=160
Y=160/2
Y=80 POUNDS OF THE $9.20 TEA IS USED.
X+80=100
X=100-80
X=20 POUNDS OF THW $7.20 TEA IS USED.
PROOF:
7.20*20+9.20*80=880
144+736=880
880=880